Combining 0.381 mol Fe2O3 with excess carbon produced 18.3 g Fe. Fe2O3+3C⟶2Fe+3CO What is the actual yield of iron in moles? actual yield: mol What is the theoretical yield of iron in moles? theoretical yield: mol What is the percent yield? percent yield:
Answer #1

Fe2O3 + 3 C ----------------- 2 Fe + 3 CO

1 mole                                 2 mole

a)

mass of Fe produced = 18.3 grams

molar mass of Fe = 55.85 gram/mole

number of moles of Fe formed = mass / molar mass = 18.3 / 55.85 = 0.328 moles

Actual yield of Iron in number of moles = 0.328 moles

b)

Given number of moles of Fe2O3 = 0.381 mole

according to balanced equation

1 mole of Fe2O3 = 2 moles of Fe

0.381 moles of Fe2O3 = ?

                                      = 2 x 0.381/1 = 0.762 moles

Theoritical yield of Fe in number of moles = 0.762 moles

c) Actual yiled = 0.328 moles

Theoritical yield = 0.762 moles

Percent yield = Actual yield x 100/Theoritical yield = 0.328 x 100/0.762 = 43.04 %

Percent yield = 43.04%.